# Beyond the Laplacian: Interpolated Spectral Augmentation for Graph Neural Networks

### v1

@@ 3.3 Richness of Representation
In defining the matrix family ${\mathbf{M}}(t,s)$, we have dropped two components corresponding to $\mathbf{I}$} and $\mathbf{J}$} in the universal adjacency matrix family ${\mathbf{M}}(\alpha,\beta,\kappa,\zeta)$ for a more parsimonious representation that is conducive to ML practice. Below, we show that dropping the $\mathbf{I}$} component does not affect the richess of our spectral embeddings. This is exemplified by the following lemma:
@@
Note that the same statement applies when $\mathbf{M}$} has repeated eigenvalues, except it would be the eigenspaces rather than then eigenvectors that are identical.
The above implies that not only are the eigenvectors of $\mathbf{M}$} unchanged when it is perturbed by $\zeta{\mathbf{I}}$, the ordering of the eigenvalues are unchanged either. In our context of selecting the first $k$ eigenvectors for spectral embeddings, the above lemma shows that ${\mathbf{M}}(t,s)$ yields the same expressivity of spectral embeddings as ${\mathbf{M}}(t,s)+\zeta{\mathbf{I}}$.
@@
${\mathbf{L}}_{\mathrm{deformed}}(q)\;=\;{\mathbf{I}}-q{\mathbf{A}}+q^{2}({\mathbf{D}}-{\mathbf{I}}),\quad q\in\mathbb{R},$ (12)
Note that in the original formulation $q\in\mathbb{C}$ is allowed for full mathematical generality, here we only consider real $q$ for our ML context. This family of matrices is important in graph theory, as it is known to capture meaningful centrality measures via non-backtracking random walks (grindrod2018deformed).
@@ 4.1 Experimental Setup
We evaluate the effectiveness of ILE in the context of node classification, where ILE is used to augment node features in GNNs. The resulting models are then assessed based on their classification accuracy. Our experiments cover a range of datasets and models.
@@
    We first consider a synthetic scenario with Stochastic Block Models (SBM) consisting of two blocks of equal size. Node labels correspond to block membership. Following priebe2019two, we choose connectivity parameters that lead to two graphs exhibiting core–periphery and community structures respectively. See Figure 2 for illustration.
@@
    We consider network datasets that do not come with node features. This include the Karate Club (karateclub), Twitter Congress (fink2023twitter), Facebook Ego (NIPS2012_7a614fd0), and Political Blogs (Polblogs) (10.1145/1134271.1134277) datasets. We use ILEs as node features on these datasets.
@@
We consider three representative models: GCN kipf2017semisupervisedclassificationgraphconvolutional, GIN xu2019powerfulgraphneuralnetworks, and GraphSAGE hamilton2018inductiverepresentationlearninglarge. We also use a 5-layer MLP as a baseline. Detailed descriptions of all datasets are provided in Appendix B.
@@ 4.2 SBM Simulations
Figure 2: Illustration of two realizations of SBM graphs that exhibit (a) core-periphery and (b) community structures
As shown in Table 1, the best-performing embeddings across the datasets are members of ILEs that do not correspond to Laplacian embeddings. Observe that the GCN model has the best performance, whereas the MLP model, which did not leverage graph convolutional operations, performed roughly at $50$ percent accuracy. Since the Core-Periphery case is largely determined by the degree representation, we would expect that the embeddings that are most informative about high degree nodes ($t=1$) or low degree nodes (t=1)t=-1)) to enjoy good performance. Indeed, for GCN, GIN and GraphSAGE, the best performance came from the t=1t=-1 and $t=1$ columns.
In the community structure case, we would expect that embeddings that are close to the Laplacian or Signless Laplacian embeddings would capture the community structure much better than adjacency based embeddings. Indeed, across the models the ILE embeddings performed better than the adjacency embeddings. Note that GCN is a special case since its convolutional operation includes a normalized Laplacian component, so it is unsurprising that it is able to detect community patterns better than other models.
@@ 4.3 Datasets without Features
In our experiments with the datasets that did not come with node features, we again consistently see the pattern where spectral feature augmentation with ILEs led to better results than the Laplacian embeddings ($s=1,t=1$). Results are shown in Table 2 in the appendix.
@@ 4.4 Datasets with Node Features Corrupted by Noise
We also considered three datasets that come with node features (Texas, Cornell and Wisconsin). Here we corrupt the original node features with Gaussian noise of different magnitude. Results for the datasets Texas, Cornell and Wisconsin are shown in Tables 5, 3 and 4, respectively, in the appendix. We observe a pattern where the larger the noise perturbation, generally the worse the performance. We also observe that at low noise corruption levels, the ILE families that perform well in general (e.g. the $t=0.5$ column in the Wisconsin dataset under 0 percent and $10$ percent corruption in Table 4) are generally stable. At high noise corruption level, the numerous noisy node features appear to overpower the effect of the spectral embeddings.
@@ 4.5 Selecting Hyperparameters
A main goal of this paper is to show that more general spectral embeddings can outperform Laplacian embeddings. While this observation holds true in our experiments which look at options in an exhaustive fashion, in practice users might have to select certain tuning parameters, including the dimension of the spectral embeddings $k$, as well as the tuning parameters $t,s$ for the interpolated Laplacian family.
For selecting the spectral dimension $k$ for Laplacian-based families, the multi-way Cheeger’s inequality lee2014multiway suggests that $k$ should be selected as the number of communities in the underlying graph. If no such prior information is available, then a conservative $k$ (an upper bound on the number of communities) should be chosen. For general graph matrices, one can adopt the commonly used strategy of inspecting Scree plots and determining $k$ according to the location of the elbow (cattell1966scree).
As for the selection of $s,t$, if ones has prior information on the graph’s underlying structure (e.g. community versus core-periphery) and their relationship to the node labels, one can selectively try appropriate ranges of $s,t$ according to the interpretation provided in Section 3. A slightly more computationally intensive approach would be to compute some correlation measure between the spectral embeddings and the node labels in a training or validation set. Yet another more computationally intensive approach is to perform cross-validation to select the $s,t$ that leads to the highest validation accuracy.
@@ 4.6 Computational Time
The computational complexity of a naïve spectral decomposition of a $n\times n$ matrix is on the order of $O(n^{3})$. However, fast numerical and/or iterative schemes (e.g. Lanczos algorithm, Power method, Preconditioned methods etc) are often available and can be highly effective in practice for the computational of only the top $k$ eigenvectors (press2007numerical). Some of these iterative methods enjoy linear converge rates. Moreover, fast numerical methods are often available for sparse graphs or for the approximate computational of eigenvectors, which may suffice for the application at hand.
@@ 5 Discussion
This work broadens the scope of spectral augmentation in graph neural networks by moving beyond the Laplacian to a unified family of graph matrices that we term interpolated Laplacian Embeddings (ILEs). Our theoretical analysis provides an interpretable characterization of the structural properties that different parameterizations of ILEs capture, ranging from community to core–periphery structure. Empirically, we demonstrated that the choice of graph matrix can substantially influence downstream node classification performance, particularly in the limited- or missing-feature regime. Natural directions of future research include generalization of such approaches to directed graphs, or to graphs with edge features, and to graphs with dynamic, temporal, or heterophilic structures. The theoretical and empirical aspects of non-Laplacian operators in graph machine learning in general is understudied compared to the canonical Laplacian case. We hope that this work motivates more systematic study of spectral embeddings beyond the Laplacian, and inspires new methods that more effectively leverage different types of graph structures for representation learning.
@@ Appendix B Dataset Details
We next evaluate ILE on real-world networks that do not contain node features. In these settings, we demonstrate that augmenting the nodes with ILE substantially improves classification accuracy. The datasets in this group are Zachary’s Karate Club (karateclub), a social network of 34 nodes and 78 edges partitioned into four classes; Twitter Congress (fink2023twitter), a network of 475 nodes and 13,289 edges representing interactions among members of the 117th United States Congress; Facebook Ego (NIPS2012_7a614fd0), a social network of 4,039 nodes and 88,234 edges capturing ego-centric friendship circles on Facebook; and Political Blogs (Polblogs) (10.1145/1134271.1134277), a network of 1,490 nodes and 19,025 edges, with binary labels indicating the political leaning of online blogs. The Twitter Congress and Facebook Ego datasets lack ground-truth node labels. To enable evaluation, we construct artificial labels based on node degree: nodes in the top 20% by degree are assigned label 1, and all others are assigned label 0.
Finally, we evaluate the ILE family on real-world datasets with node attributes. In this setting, we show that augmenting node features with ILE improves classification accuracy both under standard conditions and in scenarios where the original features are corrupted. The datasets in this group are Cornell (Pei2020Geom-GCN:), a web network with 183 nodes, 298 edges, 1,703 features, and five classes, where nodes correspond to university web pages and edges represent hyperlinks; Texas Pei2020Geom-GCN:, a structurally similar university web network with 183 nodes, 325 edges, 1,703 features, and five classes; and Wisconsin (Pei2020Geom-GCN:), another university web network with 251 nodes, 515 edges, 1,703 features, and five classes.

### v2

@@ 3.3 Richness of Representation
In defining the matrix family ${\mathbf{M}}(t,s)$, we have dropped two components corresponding to $\mathbf{I}$} and $\mathbf{J}$} in the universal adjacency matrix family ${\mathbf{M}}(\alpha,\beta,\kappa,\zeta)$ for a more parsimonious representation that is conducive to machine learning practice. Below, we show that dropping the $\mathbf{I}$} component does not affect the richness of our spectral embeddings. This is exemplified by the following lemma:
@@
Note that the same statement applies when $\mathbf{M}$} has repeated eigenvalues, except it would be the eigenspaces rather than the eigenvectors that are identical.
The above implies that not only are the eigenvectors of $\mathbf{M}$} unchanged when it is perturbed by $\zeta{\mathbf{I}}$, but the ordering of the eigenvalues is also unchanged. In our context of selecting the first $k$ eigenvectors for spectral embeddings, the above lemma shows that ${\mathbf{M}}(t,s)$ yields the same expressivity of spectral embeddings as ${\mathbf{M}}(t,s)+\zeta{\mathbf{I}}$.
@@
${\mathbf{L}}_{\mathrm{deformed}}(q)\;=\;{\mathbf{I}}-q{\mathbf{A}}+q^{2}({\mathbf{D}}-{\mathbf{I}}),\quad q\in\mathbb{R}.$ (12)
Note that in the original formulation, $q\in\mathbb{C}$ is allowed for full mathematical generality; here we only consider real $q$ for our machine learning context. This family of matrices is important in graph theory, as it is known to capture meaningful centrality measures via non-backtracking random walks (grindrod2018deformed).
@@ 4.1 Experimental Setup
We evaluate the effectiveness of ILEs in the context of node classification, where ILEs are used to augment node features in GNNs. The resulting models are then assessed based on their classification accuracy. Our experiments cover a range of datasets and models.
@@
    We first consider a synthetic scenario with Stochastic Block Models (SBMs) consisting of two blocks of equal size. Node labels correspond to block membership. Following priebe2019two, we choose connectivity parameters that lead to two graphs exhibiting core–periphery and community structures, respectively. See Figure 2 for illustration.
@@
    We consider network datasets that do not come with node features. These include the Karate Club (karateclub), Twitter Congress (fink2023twitter), Facebook Ego (NIPS2012_7a614fd0), and Political Blogs (Polblogs) (10.1145/1134271.1134277) datasets. We use ILEs as node features on these datasets.
@@
We consider three representative models: GCN (kipf2017semisupervisedclassificationgraphconvolutional), GIN (xu2019powerfulgraphneuralnetworks), and GraphSAGE (hamilton2018inductiverepresentationlearninglarge). We also use a 5-layer MLP as a baseline. Detailed descriptions of all datasets are provided in Appendix B.
@@ 4.2 SBM Simulations
Figure 2: Illustration of two realizations of SBM graphs that exhibit (a) core-periphery and (b) community structures.
As shown in Table 1, the best-performing embeddings across the datasets are members of ILEs that do not correspond to Laplacian embeddings. Observe that the GCN model has the best performance, whereas the MLP model, which did not leverage graph convolutional operations, performed roughly at $50$ percent accuracy. Since the Core-Periphery case is largely determined by the degree representation, we would expect that the embeddings that are most informative about high-degree nodes ($t=1$) or low-degree nodes (t=1t=-1) to enjoy good performance. Indeed, for GCN, GIN, and GraphSAGE, the best performance came from the t=1t=-1 and $t=1$ columns.
In the community structure case, we would expect that embeddings that are close to the Laplacian or Signless Laplacian embeddings would capture the community structure much better than adjacency-based embeddings. Indeed, across the models, the ILE embeddings performed better than the adjacency embeddings. Note that GCN is a special case since its convolutional operation includes a normalized Laplacian component, so it is unsurprising that it is able to detect community patterns better than other models.
@@ 4.3 Datasets without Features
In our experiments with the datasets that did not come with node features, we again consistently see the pattern where spectral feature augmentation with ILEs leads to better results than the Laplacian embeddings ($s=1,t=1$). Results are shown in Table 2 in the appendix.
@@ 4.4 Datasets with Node Features Corrupted by Noise
We also considered three datasets that come with node features (Texas, Cornell, and Wisconsin). Here, we corrupt the original node features with Gaussian noise of different magnitudes. Results for the datasets Texas, Cornell, and Wisconsin are shown in Tables 5, 3, and 4, respectively, in the appendix. We observe a pattern where the larger the noise perturbation, generally the worse the performance. We also observe that at low noise corruption levels, the ILE families that perform well in general (e.g. the $t=0.5$ column in the Wisconsin dataset under 0 percent and $10$ percent corruption in Table 4) are generally stable. At high noise corruption levels, the numerous noisy node features appear to overpower the effect of the spectral embeddings.
@@ 4.5 Selecting Hyperparameters
A main goal of this paper is to show that more general spectral embeddings can outperform Laplacian embeddings. While this observation holds true in our experiments, which look at options in an exhaustive fashion, in practice users might have to select certain tuning parameters, including the dimension of the spectral embeddings $k$, as well as the tuning parameters $s,t$ for the interpolated Laplacian family.
For selecting the spectral dimension $k$ for Laplacian-based families, the multi-way Cheeger’s inequality (lee2014multiway) suggests that $k$ should be selected as the number of communities in the underlying graph. If no such prior information is available, then a conservative $k$ (an upper bound on the number of communities) should be chosen. For general graph matrices, one can adopt the commonly used strategy of inspecting Scree plots and determining $k$ according to the location of the elbow (cattell1966scree).
As for the selection of $s,t$, if one has prior information on the graph’s underlying structure (e.g. community versus core-periphery) and their relationship to the node labels, one can selectively try appropriate ranges of $s,t$ according to the interpretation provided in Section 3. A slightly more computationally intensive approach would be to compute some correlation measure between the spectral embeddings and the node labels in a training or validation set. Yet another more computationally intensive approach is to perform cross-validation to select the $s,t$ that leads to the highest validation accuracy.
@@ 4.6 Computational Time
The computational complexity of a naïve spectral decomposition of an $n\times n$ matrix is on the order of $O(n^{3})$. However, fast numerical and/or iterative schemes (e.g. Lanczos algorithm, Power method, Preconditioned methods, etc.) are often available and can be highly effective in practice for the computation of only the top $k$ eigenvectors (press2007numerical). Some of these iterative methods enjoy linear convergence rates. Moreover, fast numerical methods are often available for sparse graphs or for the approximate computation of eigenvectors, which may suffice for the application at hand.
@@ 5 Discussion
This work broadens the scope of spectral augmentation in GNNs by moving beyond the Laplacian to a unified family of graph matrices that we term Interpolated Laplacian Embeddings (ILEs). Our theoretical analysis provides an interpretable characterization of the structural properties that different parameterizations of ILEs capture, ranging from community to core–periphery structure. Empirically, we demonstrated that the choice of graph matrix can substantially influence downstream node classification performance, particularly in the limited- or missing-feature regimes. Natural directions of future research include generalization of such approaches to directed graphs, to graphs with edge features, and to graphs with dynamic, temporal, or heterophilic structures. The theoretical and empirical aspects of non-Laplacian operators in graph machine learning in general are understudied compared to the canonical Laplacian case. We hope that this work motivates more systematic studies of spectral embeddings beyond the Laplacian and inspires new methods that more effectively leverage different types of graph structures for representation learning.
@@ Appendix B Dataset Details
We next evaluate ILEs on real-world networks that do not contain node features. In these settings, we demonstrate that augmenting the nodes with ILEs substantially improves classification accuracy. The datasets in this group are Zachary’s Karate Club (karateclub), a social network of 34 nodes and 78 edges partitioned into four classes; Twitter Congress (fink2023twitter), a network of 475 nodes and 13,289 edges representing interactions among members of the 117th United States Congress; Facebook Ego (NIPS2012_7a614fd0), a social network of 4,039 nodes and 88,234 edges capturing ego-centric friendship circles on Facebook; and Political Blogs (Polblogs) (10.1145/1134271.1134277), a network of 1,490 nodes and 19,025 edges, with binary labels indicating the political leaning of online blogs. The Twitter Congress and Facebook Ego datasets lack ground-truth node labels. To enable evaluation, we construct artificial labels based on node degree: nodes in the top 20% by degree are assigned label 1, and all others are assigned label 0.
Finally, we evaluate the ILE family on real-world datasets with node attributes. In this setting, we show that augmenting node features with ILEs improves classification accuracy both under standard conditions and in scenarios where the original features are corrupted. The datasets in this group are Cornell (Pei2020Geom-GCN:), a web network with 183 nodes, 298 edges, 1,703 features, and five classes, where nodes correspond to university web pages and edges represent hyperlinks; Texas Pei2020Geom-GCN:, a structurally similar university web network with 183 nodes, 325 edges, 1,703 features, and five classes; and Wisconsin (Pei2020Geom-GCN:), another university web network with 251 nodes, 515 edges, 1,703 features, and five classes.